### v3

# On Conjecture of Binomial Edge Ideals of Linear Type

Abstract.
An ideal $I$ of a commutative ring $R$ is said to be of linear type when its Rees algebra and symmetric algebra exhibit isomorphism. In this paper, we investigate the conjecture put forth by Jayanthan, Kumar, and Sarkar in [11] that if $G$ is a tree or a unicyclic graph, then the binomial edge ideal of $G$ is of linear type. Our investigation validates this conjecture for trees. However, our study reveals that not all unicyclic graphs adhere to this conjecture.
Key words and phrases:
Rees algebra, linear type, $d$-sequence, $p$-sequence, binomial edge ideal, tree.
2020 Mathematics Subject Classification:
Primary 05E40, 13F65, 13F70; Secondary 13A30
Introduction
Consider a graph $G$ on $[n]$ vertices. Let $S=\mathbb{K}[x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}]$ be a polynomial over a field $\mathbb{K}$. Herzog et al. in [7] and independently Ohtani in [15] introduced the ideal
$J_{G}=(f_{ij}=x_{i}y_{j}-x_{j}y_{i}\mid i<j,\{i,j\}\in E(G))\subset S$
is called binomial edge ideal of the graph $G$. We call the binomial $f_{ij}$ associated to an edge $\{i,j\}\in E(G)$ as an edge binomial. In [7], Herzog et al. observed that the binomial edge ideal appears naturally in the study of conditional independence ideals, providing a valuable tool to investigate robustness theory in the context of algebraic statistics. In recent years, researchers have investigated various algebraic properties, including radicality, primality, Gröbner bases, primary decomposition, among others, of binomial edge ideals in relation to their underlying combinatorial properties; see for partial list [5, 7, 14, 15, 18] and the references therein.
Let $A$ be a standard graded polynomial ring over a field $\mathbb{K}$ and $I\subset A$ be a homogenous ideal. The Rees algebra of an ideal $I\subset A$, is defined as $\mathcal{R}(I)=\{\sum_{i=0}^{n}a_{i}t^{i}\mid n\in\mathbb{N}$ and $a_{i}\in I^{i}\}=\bigoplus_{i\geq 0}I^{i}t^{i}\subset A[t]$, where $I^{0}=A$. $\mathcal{R}(I)$ is a classical algebraic structure that encodes numerous asymptotic properties of that ideal. The symmetric algebra of an ideal $I\subset A$ is defined as $Sym(I)=T(I)/L$, where $T(I)$ is the tensor algebra of $I$ and $L$ is the ideal of $T(I)$ generated by elements of the form $x\otimes_{A}y-y\otimes_{A}x$, where $x,y\in I$. Moreover, there exists a natural epimorphism from Sym$(I)$ to $\mathcal{R}(I)$, where Sym$(I)$ is the symmetric algebra of $I$. For an ideal $I\subset A$, if Sym$(I)\cong\mathcal{R}(I)$, then $I$ is referred to as an ideal of linear type. The defining equations of the symmetric algebra of an ideal can always be computed from a presentation matrix of the ideal.
Villareal provided a characterization of the ideal of linear type for edge ideals, namely, edge ideals are of linear type if and only if the graph is a tree or an odd unicyclic graph (cf. [21]). This raises the question of whether a similar characterization might apply to other combinatorial objects. In [11], Jayanthan et al. initiated such a study for binomial edge ideals. In their work, the authors characterized almost complete intersection ideals and proposed the following conjecture:
Conjecture A.
[11, Conjecture 4.17. (a)] If $G$ is a tree or a unicyclic graph, then $J_{G}$ is of linear type.
Huneke introduced the notion of $d$-sequence in [9, 10]. Huneke, and independently Valla, proved that if the ideal is generated by unconditioned $d$-sequence, then the ideal is of linear type in [9, 20]. Costa [4] established that $d$-sequence condition on $I$ is a sufficient condition for the ideal $I$ to be of linear type. Later, Raghavan [16] provided a simpler proof that $d$-sequence ideals are of linear type. To the best of our knowledge, study of the Rees algebra of binomial edge ideal of linear type are the following. In [11], the authors characterised almost complete intersection binomial edge ideals, a subclass of ideal of linear type. In [1, 2], the authors provided a partial answer for Conjecture A. Particularly, the authors characterized $d$-sequence binomial edge ideals of trees and unicyclic graphs at the level of a sequence of edge binomials. In [13], Kumar characterized binomial edge ideals of linear type for closed graphs.
In this paper, we settle Conjecture A for trees; see Theorem 4.1. To achieve this, we introduce a $p$-sequence in Section 2, as defined in Definition 2.1. In Section 3, we proved that trees and a class of unicyclic graphs form a $p$-sequence, see Theorems 3.1 and 3.2. Finally, in Section 5, we present a counterexample illustrating that the conjecture does not hold for unicyclic graphs in general; refer to Proposition 5.1.
1. Preliminaries
In this section, we recall basic definitions, notations, and results from the literature that are utilized in the subsequent sections.
1.1. Basic notions from graph theory.
Let $G=(V(G)=[n],E(G))$ be a finite simple graph. For a subset $U\subseteq V(G)$, $G[U]$ denotes induced subgraph on $U$ of $G$, defined as $G[U]=(U,\{\{i,j\}\in E(G)\mid i,j\in U\})$. A graph $G$ is said to be path if $E(G)=\{\{i,i+1\}\mid 1\leq i\leq n-1\}\}$, and it is denoted by $P_{n}$. A graph $G$ is said to be cycle if $E(G)=\{\{1,n\}\cup\{i,i+1\}\mid 1\leq i\leq n-1\}\}$, for $n\geq 3$, and it is denoted by $C_{n}$. A graph $G$ is said to be forest if it does not have a cycle and a tree if it is connected. It is said to be unicyclic graph if it contains precisely one cycle as a subgraph.
Notation 1.1.
Let $G=(V(G),E(G))$ be a graph. For an edge $e$ and a vertex $v$ in $G$:
  • $G\setminus e$ is the graph on the vertex set $V(G)$ and edge set $E(G)\setminus{e}$.
  • $e$ is called a bridge if $c(G)<c(G\setminus e)$, where $c(G)$ is the number of components of $G$.
  • $N_{G}(v)=\{u\in V(G)\mid\{u,v\}\in E(G)\}$ denotes the neighborhood of $v$ in $G$.
  • The degree of a vertex $v\in V(G)$, denoted by $\deg_{G}(v)$, is defined as the number of edges incident to $v$. If $\deg_{G}(v)=1$, then $v$ is called a pendant vertex.
  • An edge $e$ is called a pendant edge if one of the vertices is a pendant vertex.
  • [14, Definition 3.1] Let $e^{\prime}=\{i,j\}\notin E(G)$ be an edge in $G\cup\{e^{\prime}\}$. Then $G_{(e^{\prime})}$ is the graph on vertex set $V(G)$ and edge set
    $E(G_{(e^{\prime})})=E(G)\cup\left\{\{k,l\}\mid k,l\in N_{G}(i)\textnormal{ or }k,l\in N_{G}(j)\right\}.$
A rooted tree $(G,x_{0})$ is a directed tree in which edges are implicitly directed away from the root pendant vertex $x_{0}$. A level of a vertex $u\in V(G)$ in a rooted tree $(G,x_{0})$ is the length of the path $P$ from $x_{0}$ to $u$. A level of an edge $e=\{x_{i},x_{j}\}$ in a rooted tree $(G,x_{0})$ is the length of the path $P$ from $x_{0}$ to $x_{j}$ such that $x_{i}\in V(P)$.
1.2. Basic results from commutative algebra.
Let $G$ denote a finite simple graph on vertex set $[n]$. Define $S=\mathbb{K}[x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}]$ as a polynomial ring over the field $\mathbb{K}$, and let $J_{G}$ be an ideal generated by edge binomials that are derived from $G$. Throughout this article, the symbols $S$ and $J_{G}$ are exclusively used to refer to the polynomial ring and the binomial edge ideal, respectively. The term “sequence of elements in a ring” refers to a list of elements within that ring.
The subsequent results describe the colon ideal operation of an edge binomial on the binomial edge ideal, which we use repeatedly throughout this article.
Remark 1.1.
[11, Lemma 4.1] [7, Corollary 2.2] Let $G$ be a simple graph and $f$ be an element of $S$. Then $J_{G}:f=J_{G}:f^{n}$ for all $n\geq 2$.
Remark 1.2.
[14, Theorem 3.4] Let $G$ be a simple graph. Let $e=\{i,j\}\notin E(G)$ be a bridge in $G\cup\{e\}$. Then $J_{G}:f_{e}=J_{G_{e}}$.
Remark 1.3.
[14, Theorem 3.7] Let $G$ be a simple graph and $e=\{i,j\}\notin E(G)$ be an edge. Then
$J_{G}:f_{e}=J_{G_{e}}+(g_{P,t}\mid\quad P:i,i_{1},\ldots,i_{s},j\text{ is a path between }i,j\text{ and }0\leq t\leq s),$
where $g_{P,0}=x_{i_{1}}\ldots x_{i_{s}}$ and for each $1\leq t\leq s,g_{P,t}=y_{i_{1}}\ldots y_{i_{t}}x_{i_{t+1}}\ldots x_{i_{s}}$.
Remark 1.4.
[18, Corollary 2.5] For any bipartite graph $G$, one has $\beta_{i,i+1}(S/J_{G})=0$, for all $i$.
Definition 1.1.
[10, Definition 1.1] Let $A$ be a commutative ring. Set $z_{0}=0$. A sequence of elements $z_{1},\ldots,z_{m}$ in $A$ is said to be a $d$-sequence if it satisfies the conditions:
  1. (1)
    $z_{1},\ldots,z_{n}$ is a minimal system of generators of the ideal $(z_{1},\ldots,z_{n})$;
  2. (2)
    $((z_{0},\ldots,z_{i}):z_{i+1}z_{j})=((z_{0},\ldots,z_{i}):z_{j})$ for all $10\leq i\leq n-$ 1, and $j\geq i+1$.
If $z_{1},\ldots,z_{m}$ is a $d$-sequences in any order then the sequence is said to be unconditional $d$-sequence.
Definition 1.2.
Let $A$ be a polynomial ring over a field $\mathbb{K}$.
  1. (a)
    An ideal $I\subset A$ is said to be complete intersection if $\mu(I)=ht(I)$, where $\mu(I)$ denotes the cardinality of a minimal homogeneous generating set of $I$.
  2. (b)
    An ideal $I\subset A$ is said to be an almost complete intersection if $\mu(I)=\operatorname{ht}(I)+1$ and $I_{p}$ is a complete intersection for all minimal primes $p$ of $I$.
Remark 1.5.
[11, Proposition 4.10] Let $A$ be a polynomial ring over a field $\mathbb{K}$. Let $I\subset A$ be a homogeneous almost complete intersection ideal. Then $I$ is generated by a homogeneous $d$-sequence $z_{1},\ldots,z_{h+1}$ such that $z_{1},\ldots,z_{h}$ is a regular sequence, $h=\operatorname{ht}(I)$.
Definition 1.3.
[17, Definition 1.1] A relation on an ordered sequence of elements $z_{1},\ldots,z_{n}$ of a commutative ring $A$ is a form of positive degree in the graded polynomial ring $R=A[X_{i}\mid 1\leq i\leq n]$, where elements of $A$ have degree 00 and each $X_{i}$ has degree $1$, that yields 00 when evaluated at $z_{1},\ldots,z_{n}$. Let $\Lambda_{j}$ denote the ideal of $R$ generated by the relations on $z_{1},\ldots,z_{n}$ of degree at most $j$. Let $r$ be the least positive integer such that $\Lambda_{r}=\cup_{j\geq r}\Lambda_{j}$. The invariant $r$ is called the relation type of the ideal $I=(z_{1},\ldots,z_{n})$. An ideal of relation type $1$ is said to be of linear type.
Throughout this paper, we reserve $R$ for the polynomial ring over $A$ unless otherwise specified.
Definition 1.4.
[4, p. 257] Let $A$ be a commutative ring. An ordered sequence of elements $z_{1},\ldots,z_{n}$ of a ring $A$ is said to be a sequence of linear type if each of the ideal $(z_{1},\ldots,z_{i})$ is of linear type for all $1\leq i\leq n$.
2. p-sequence
In this section, we introduce a new notion called $p$-sequence.
Definition 2.1.
Let $A$ be a commutative ring. Set $z_{0}=0$. A sequence of elements $z_{1},\ldots,z_{n}$ in $A$ is said to be $p$-sequence if it satisfies the conditions:
  1. (1)
    $z_{1},\ldots,z_{n}$ is a minimal system of generators of the ideal $(z_{1},\ldots,z_{n})$;
  2. (2)
    $(z_{0},\ldots,z_{i}):z_{i_{1}}z_{i_{2}}=(z_{0},\ldots,z_{i}):z_{i_{1}}$ for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq i_{2}\leq n$.
An ideal is said to be a $p$-sequence if a minimal generating set of the ideal forms a $p$-sequence.
Clearly, unconditional $d$-sequences are $p$-sequences. Examples of a class of unconditional $d$-sequences were given in [10]. We list some examples of $p$-sequences here.
Example 1.
Any regular sequence is a $p$-sequence.
Proposition 2.1.
Let $A$ be a commutative ring. Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $A$ forms a $p$-sequence. Then $z_{i+1},\ldots,z_{n}$ in $A/I_{i}$ forms a $p$-sequence, where $I_{i}=(z_{1},\ldots,z_{i})$, for all $i$.
Proof.
It follows from Definition 2.1. ∎
In the subsequent examples, we show that up to the sequence level, $d$-sequences and $p$-sequences are distinct notions.
Example 2.
Let $G$ be a graph with vertex set $V(G)=[8]$ and edge set
$E(G)=\{\{1,2\},\{1,3\},\{1,4\},\{1,5\},\{5,6\},\{5,7\},\{5,8\}\}.$
Then from Theorem 3.1, it follows that there exists a sequence of edge binomials of $G$ that form a $p$-sequence. However, the edge binomials of $G$ does not form a $d$-sequence under any permutation . This is due to the fact that for any graph $G$ having at least two vertices with degrees greater than or equal to $4$, edge binomials of $G$ can not form a $d$-sequence (cf. [1, Corollary 3.2]).
Example 3.
Consider a sequence of elements $\{x_{1}x_{3}x_{4}x_{5},x_{1}^{2}x_{2}x_{6},x_{1}^{2}x_{2}^{2}x_{3}x_{5}\}$ in $\mathbb{K}[x_{1},\ldots,x_{6}]$. In the specified order, this sequence readily forms a $d$-sequence. However, no permutation of this sequence constitutes a $p$-sequence. This assertion is verified by Computation 5.1.
Definition 2.2.
Let $A=\mathbb{K}[x_{1},\ldots,x_{n}]$ be a polynomial ring over a field $\mathbb{K}$. Any product of variables of the form $x^{\alpha}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$ is called monomial, where $\alpha\in\mathbb{N}^{n}$.
Proposition 2.2.
Let $z_{1},\ldots,z_{n}$ be a monomial sequence in $\mathbb{K}[x_{1},\ldots,x_{n}]$. Then $z_{1},\ldots,z_{n}$ is a $p$-sequence if for $i\neq j$ one has $z_{i}\nmid z_{j}$ and
$\gcd(z_{i},z_{i_{2}})|z_{i_{1}},1\leq i<i_{1}\leq i_{2}\leq n,$ (2.1)
$\gcd(z_{i},z_{j})|\gcd(z_{i},z_{j}^{2}),1\leq i<j\leq n.$ (2.2)
Proof.
To demonstrate that $z_{1},\ldots,z_{n}$ forms a $p$-sequence, it is enough to show that for any $i<i_{1}\leq i_{2}$, $\gcd(z_{i},z_{i_{1}}z_{i_{2}})=\gcd(z_{i},z_{i_{1}})$. From Equation (2.2) it follows that $\gcd(z_{i},z_{i_{1}}^{2})=\gcd(z_{i},z_{i_{1}})$ for all $i_{1}>i$. From Equation (2.1) it follows that $\gcd(z_{i},z_{i_{1}}z_{i_{2}})\mid\gcd(z_{i},z_{i_{1}}^{2})$. Then by Equation (2.2) one has $\gcd(z_{i},z_{i_{1}}z_{i_{2}})=\gcd(z_{i},z_{i_{1}})$, for all $i<i_{1}<i_{2}$, as desired.
Remark 2.1.
In the context of edge ideals, Proposition 2.2 implies that edge ideals of matching or caterpillar graphs form a $p$-sequence. These graphs evidently belong to the subset of ideals characterized as ideals of linear type by Villareal [21]. Thus, it prompts the natural question.
Question 2.1.
Let $A$ be a commutative ring. Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $A$ form a $p$-sequence. Is the ideal $I=(z_{1},\ldots,z_{n})$ of linear type?
The following proposition facilitates addressing the above question, as it allows consideration of only relation polynomials where each monomial is squarefree.
Proposition 2.3.
Let $A$ be a standard graded polynomial ring over an field. Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $A$ forms a $p$-sequence such that it satisfy
$z_{i}\cap(z_{1},\ldots,z_{i-1})^{s}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n})\subseteq z_{i}(z_{1},\ldots,z_{i-1})^{s-1}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n})$ (2.3)
for all $i=1,\ldots,n$ and $s\geq 1$. Then for any $F(X_{1},\ldots,X_{n})$ a homogeneous positive degree polynomial such that $F(z_{1},\ldots,z_{n})=0$, can be linearized in $X_{1},\ldots,X_{n}$.
Proof.
The proof proceeds by induction on $\deg(F)$. If $\deg(F)=1$, there is nothing to prove. Suppose $\deg(F)=d$ and $F(z_{1},\ldots,z_{n})=0$, such that $F$ is linear in $X_{n},\ldots,X_{i+1}$. Rewrite $F(X_{1},\ldots,X_{n})$ as $X_{i}G(X_{1},\ldots,X_{n})+H(X_{1},\ldots,\hat{X_{i}},\ldots,X_{n})$, where $\deg(G)=d-1$, $\deg(H)=d$, and both $G$ and $H$ are linear in $X_{n},\ldots,X_{i+1}$. Let $\bar{H}$ denote the polynomial $H(z_{1},\ldots,z_{n})$. It follows that $\bar{H}\in(z_{i})$, since $F(z_{1},\ldots,z_{n})=0$. This shows that
$\bar{H}\in(z_{i})\cap(z_{1},\ldots,z_{i-1})^{d-1}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n}).$
From Equation (2.3) it follows that $\bar{H}\in z_{i}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n})(z_{1},\ldots,z_{i-1})^{d-2}$. Thus $\bar{H}$ can be written as $\bar{H}=z_{i}G^{\prime}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n})$, where $G^{\prime}(X_{1},\ldots,\hat{X_{i}},\ldots,X_{n})$ is linear in $X_{i+1},\ldots,X_{n}$. This implies that $z_{i}G(z_{1},\ldots,z_{n})+z_{i}G^{\prime}(z_{1},\ldots,\hat{z_{i}},\ldots,z_{n})=0$. Therefore, one concludes that $(G+G^{\prime})(z_{1},\ldots,z_{n})\in(0:z_{i})\cap(z_{1},\ldots,z_{n})$. Given that $A$ is a domain, it follows that $(0:z_{i})=0$ for all i. Consequently, $(G+G^{\prime})(z_{1},\ldots,z_{n})=0$. Since $\deg(G+G^{\prime})<d$, it follows from the induction hypothesis that $X_{i}G+X_{i}G^{\prime}$ is linear. It remains to show that $H-X_{i}G^{\prime}$ is linear. But $H$ and $G^{\prime}$ does not have $X_{i}$ thus it follows that $H-X_{i}G^{\prime}$ is linear in $X_{n},\ldots,X_{i}$. By continuing this process, one can linearize $F$ in $X_{1},\ldots,X_{n}$. ∎
The following example demonstrates that a $p$-sequence fails to satisfy one of the properties of a $d$-sequence.
Example 4.
Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $S$ form a $d$-sequence. Then, it satisfies Equation (2.3), see [9, Proposition 3.1]. However, this property does not hold for $p$-sequences. For instance, consider a graph $G$ with edge set
$E(G)=\{\{1,2\},\{2,3\},\{2,4\},\{4,5\},\{4,6\},\{3,5\}\}.$
From Theorem 3.2, it follows that the ideal $J_{G}$ forms a $p$-sequence. Then, one can verify that
$f_{24}\cap(f_{12},f_{23})^{2}(f_{12},f_{23},f_{45},f_{46},f_{35})\not\subset f_{24}(f_{12},f_{23})(f_{12},f_{23},f_{45},f_{46},f_{35}).$
3. Binomial edge ideals of p-sequence
In this section, we demonstrate that the edge binomials of trees and a class of unicyclic graphs form a $p$-sequence.
3.1. Ordering on edges of a tree
Let $G$ be a tree. Any tree can be represented as a rooted tree structure with a pendant vertex as a root vertex, denoted by $(G,x_{0})$. The set of all edges in level $m$ is denoted by $e_{m,1},e_{m,2},\ldots,e_{m,k_{m}}$ (arranged from left to right). We ordered the edge binomials of a tree in the following order. Let $(G,x_{0})$ be a rooted tree with maximum level $m$. Then we arrange the edges by $e_{1,1},e_{2,1},\ldots,e_{2,k_{2}},\ldots,e_{m-1,1},\ldots,e_{m-1,k_{m-1}},e_{m,1},\ldots,e_{m,k_{m}}$.
Example 5.
[…]
every picture/.style=line width=0.75pt
{tikzpicture}
[x=0.75pt,y=0.75pt,yscale=-1,xscale=1]
[…]
Figure 1. The rooted tree $(G,1)$
Let $(G,1)$ be a rooted tree as given in Figure 1. Then, we arrange the edges of $(G,1)$ in the following order $e_{1,1}=\{1,2\},e_{2,1}=\{2,3\},e_{2,2}=\{2,4\},e_{2,3}=\{2,5\},e_{3,1}=\{3,6\},e_{3,2}=\{3,7\},e_{3,3}=\{3,8\},e_{3,4}=\{5,9\},e_{3,5}=\{5,10\}$.
3.2. Ordering on edges of a unicyclic graph
Let $(T,1)$ be a rooted tree with the maximal level of a vertex is $m$. Let $G$ be a unicyclic graph obtained by adding an edge $e=\{u,v\}$ between two pendant vertices of tree $T$. Represent the tree $T$ by rooted tree structure as in 3.1 such that one of the vertices of $e$ is left most in level $m$ vertex. Then we arrange the edges of $G=(T,1)\cup e$ by $e_{1,1},e_{2,1},\ldots,e_{2,k_{2}},\ldots,e_{m-1,1},\ldots,e_{m-1,k_{m-1}},e_{m,1},\ldots,e_{m,k_{m}},e$.
Example 6.
[…]
Theorem 3.1.
Let $G$ be a tree on $[n+1]$. Then, the binomial edge ideal of $G$ is a $p$-sequence ideal.
Proof.
We claim that an ordered sequence of edge binomials with respect to the edges arranged as described in Subsection 3.1 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ as ordered in Subsection 3.1 and set $d_{0}=0$. Let $e_{i}$ be the edge of $G$ associated with the edge binomial $d_{i}$ for all $i$. Now we show that
$\begin{split}(d_{0},\ldots,d_{i}):d_{i_{1}}d_{i_{2}}&=((d_{0},\ldots,d_{i}):d_{i_{1}}):d_{i_{2}}\\ &=(d_{0},\ldots,d_{i}):d_{i_{1}}\end{split}$
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq i_{2}\leq n$.
Let $H$ be the graph formed by the edges $e_{1},\ldots,e_{i}$, for some $i<n$. Notice that the graph $H$ intersect with the edge $e_{j}$ in at most one vertex, due to the arrangement of the edges of $G$, where $i\in\{0,\ldots,n-1\}$ and $j>i$. From Remark 1.2 if follows that $(d_{0},\ldots,d_{i}):d_{i_{1}}=J_{H_{(e_{i_{1}})}}$.
There are three possible way in which the edge $e_{i_{2}}=\{u,v\}$ may intersect with the graph $H$ and the edge $e_{i_{1}}$: Case I. If $e_{i_{2}}\cap H=\emptyset$, then it follows from Remark 1.2 that $J_{H_{(e_{i_{1}})}}:d_{i_{2}}=J_{H_{(e_{i_{1}})}}$. Case II. If $e_{i_{2}}\cap H\neq\emptyset$, and $e_{i_{2}}\cap e_{i_{1}}=\emptyset$, then by the arrangement of the edges of $G$ it follows that $|N_{H_{(e_{i_{1}})}\cup e_{i_{2}}}(u)|=2$ and $|N_{H_{(e_{i_{1}})}\cup e_{i_{2}}}(v)|=1$. Therefore, by applying Remark 1.2 one yields $J_{H_{(e_{i_{1}})}}:d_{i_{2}}=J_{H_{(e_{i_{1}})}}$. Case III. If $e_{i_{2}}\cap H\neq\emptyset$, and $e_{i_{2}}\cap e_{i_{1}}\neq\emptyset$, then either $e_{i_{2}}=e_{i_{1}}$ or share a common vertex, in both the situation neighborhoods of $e_{i_{2}}$ and $e_{i_{1}}$ in $H$ are equal. Thus, the equality follows from Remark 1.2. Hence, the edge binomial $d_{i_{2}}$ behaves as a non zero divisor modulo $(d_{0},\ldots,d_{i}):d_{i_{1}}$ as desired. ∎
Theorem 3.2.
Let $T$ be a tree on $n$. Suppose $G$ be a graph obtained by adding an edge between two pendant vertices of tree $T$. Then, the edge binomials of $G$ form a $p$-sequence.
Proof.
We claim that an ordered sequence of edge binomials with respect to the edges arranged as in Subsection 3.2 forms a $p$-sequence. Say $d_{1},\ldots,d_{n}$ be the ordered sequence of edge binomials of $G$ and set $d_{0}=0$. Let $e_{i}$ be the edge of $G$ associated with the edge binomial $d_{i}$ for all $i$. From Theorem 3.1 it follows that $(d_{0},\ldots,d_{i}):d_{i_{1}}d_{i_{2}}=(d_{0},\ldots,d_{i}):d_{i_{1}}$ for all $20\leq i\leq n-$ 2, and $i<i_{1}\leq i_{2}\leq n-1$. It remains to show that
$(d_{0},\ldots,d_{i}):d_{i_{1}}d_{n}=(d_{0},\ldots,d_{i}):d_{i_{1}}$
for all $10\leq i\leq n-$ 1, and $i<i_{1}\leq n$.
Observe that if $i<n$, then the edge $e_{i_{1}}=\{r,s\}$ is a leaf in the graph formed by edges $e_{1},\ldots,e_{i},e_{i_{1}}$, and the edge $e_{n}=\{u,v\}$ is a leaf in the graph formed by edges $e_{1},\ldots,e_{i},e_{n}$. Furthermore, we may assume that $s$ and $v$ are pendent vertices in their respective graphs.
For $i_{1}<n$, the equality can be obtained by applying Remark 1.2, since the edge $e_{n}$ is a leaf attached to the graph $H_{(e_{i_{1}})}$, where $H$ denotes the graph formed by the edges $e_{1},\ldots,e_{i}$. We now discuss the possible cases: Case I. If $e_{n}\cap H=\emptyset$, then the equality follows from Remark 1.2. Case II. If $e_{n}\cap H\neq\emptyset$, then there exists an edge $e_{j}$, with $j\leq i$, such that $e_{n}\cap e_{j}\neq\emptyset$. In this case, the vertex $u$ intersect with $e_{j}$. If $deg_{H_{(e_{i_{1})}}\cup e_{n}}(u)=2$ then again the equality follows from Remark 1.2. Suppose if $deg_{H_{(}e_{i_{1})}\cup e_{n}}(u)>2$. Then, it follows from the arrangement of edges of $G$ that $N_{H\cup e_{i_{1}}}(r)=N_{H_{(e_{i_{1})}}\cup e_{n}}(u)$. Now by applying Remark 1.2 one yields the equality.
For $i_{1}=n$, the equality follows from Remark 1.1. Hence, $J_{G}$ is a $p$-sequence ideal. ∎
4. On Conjecture
In this section, we aim to prove Conjecture A for all trees. We begin by establishing the necessary notations.
Notation 4.1.
Let $G=(V(G),E(G))$ be a graph with $\lvert E(G)\rvert=m$. We arrange the edge of $G$ in an ordered sequence $e_{1},\ldots,e_{m}$. We denote $G_{i}$ as the graph with vertex set $V(G_{i})=V(G)$ and edge set $E(G_{i})=\{e_{1},\ldots,e_{i}\}$, for $1\leq i\leq m$. The graph $G_{i_{(e_{i+1})}}$ is defined as in Notation 1.1.
Lemma 4.1.
Let $G$ be a tree on vertex set $[n+1]$ and $e_{1},\ldots,e_{n}$ be an ordered sequence of edges of $G$ as described in Subsection 3.1. Let $T_{i}$ be a subgraph of $G$ with vertex set $V(T_{i})=V(G)$ and edge set $E(T_{i})=E(G)\setminus E(G_{i})$ for all $i=1,\ldots,n-1$. Then one has
$J_{G_{i_{(e_{i+1})}}}:J_{T_{i}}=J_{G_{i_{(e_{i+1})}}}$ (4.1)
for all $i=1,\ldots,n-1$. Moreover, for any $0\neq f\in J_{T_{i}}$, one has $J_{G_{i_{(e_{i+1})}}}:f=J_{G_{i_{(e_{i+1})}}}$ for all $i=1,\ldots,n-1$.
Proof.
Let the edge set of $T_{i}$ be $\{e_{i+1},\ldots,e_{n}\}$ for a fixed $i$. Then one has $J_{G_{i_{(e_{i+1})}}}:J_{T_{i}}=J_{G_{i_{(e_{i+1})}}}:f_{e_{i+1}}\cap\cdots\cap J_{G_{i_{(e_{i+1})}}}:f_{e_{n}}$. Note that for any edge $e$ in $T_{i}$, $e$ is a bridge in the graph $G_{i}\cup e$. For any edge $e=\{u,v\}$ in $T_{i}$ there are three possible ways it may intersect with the graph $G_{i}$ and the edge $e_{i+1}$. Case I. If $e\cap G_{i}=\emptyset$, then it follows from Remark 1.2 that $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Case II. If $e\cap G_{i}\neq\emptyset$ and $e\cap e_{i+1}=\emptyset$, then it follows that $\deg_{G_{i}\cup e}(u)=2$ and $\deg_{G_{i}\cup e}(v)=1$ ($u$,$v$ can be interchangeable). Then, using Remark 1.2 one yields $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Case III. If $e\cap G_{i}\neq\emptyset$ and $e\cap e_{i+1}\neq\emptyset$, then either $e=e_{i+1}$ or $e$ and $e_{i+1}$ shares a common vertex, in both the situation neighborhood of $e$ and $e_{i+1}$ in $G_{i}$ are equal. Therefore, by applying Remark 1.2 one yields $J_{G_{i_{(e_{i+1})}}}:f_{e}=J_{G_{i_{(e_{i+1})}}}$. Hence Equation 4.1 holds. The containment $J_{G_{i_{(e_{i+1})}}}:f\supset J_{G_{i_{(e_{i+1})}}}$ follows from Equation 4.1. For the other side containment, suppose $a\notin J_{G_{i_{(e_{i+1})}}}$ such that $af\in J_{G_{i_{(e_{i+1})}}}$, then it follows from Remark 1.3 that there is path from $\alpha$ to $\beta$ in $T_{i}$, where $\{\alpha,\beta\}\in V(G_{i_{(e_{i+1})}})$. Based on the hypothesis that $G$ is a tree, it is not possible, as desired. ∎
Notation 4.2.
Let $z_{1},\ldots,z_{n}$ be an ordered sequence of elements of $S$. Let $\phi:R\rightarrow S$ be a map as in Definition 1.3. We denote by $I_{i}$ an ideal generated by $X_{1},\ldots,X_{i}$ and by $\bar{I}_{i}$ an ideal generated by $z_{1},\ldots,z_{i}$. Let $G(X_{1},\ldots,X_{n})$ be a polynomial in $R$. Then $\bar{G}$ denote the polynomial $G(z_{1},\ldots,z_{n})$ in $S$.
Theorem 4.1.
If $G$ be a tree on $[n+1]$, then $J_{G}$ is of linear type.
Proof.
Let $f_{e_{1}},\ldots,f_{e_{n}}$ be an ordered sequence of edge binomials for a tree as described in Subsection 3.1. According to Theorem 3.1 the sequence $f_{e_{1}},\ldots,f_{e_{n}}$ form a $p$-sequence. We propose that for any polynomial $F(X_{1},\ldots,X_{n})$ in $R$, if evaluated at $(f_{e_{1}},\ldots,f_{e_{n}})$ results in 00, then $F$ is of linear relation type. The proof is by strong induction on $\deg(F)$. If $\deg(F)=1$, there is nothing to prove. Assume that the statement holds for all $\deg(F)<d$. For $\deg(F)=d$, let $F=\sum_{\lvert\alpha\rvert\leq d}a_{\alpha}X^{\alpha}$ be a polynomial such that evaluated at $f_{e_{1}},\ldots,f_{e_{n}}$ yields 00, where $\alpha\in\mathbb{N}^{n}$. Then, we claim that $F$ is of linear type.
Set $F_{0}=F$. The proof proceeding in following steps as $k$ runs from $1$ to $n-1$. Here, we define $F_{k}$ recursively in Step $k$.
Step $k$.
Later, $F_{k-1}$ mod $I_{n-k}$ have the following cases:
Case I. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$ mod $I_{n-k}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$. Then it follows 00 mod $\bar{I}_{n-k}$, since $\bar{F}=0$. Thus $af_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}\in\bar{I}_{n-k}$. From Proposition 2.1 and Definition 2.1 it follows that $a\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Thus $af_{e_{n-(k-1)}}\in\bar{I}_{n-k}$. Therefore one has $af_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k}=aX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{k}=F_{k-1}-H_{k}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$.
Case II. If $F_{k-1}(X_{1},\ldots,X_{n})=aX_{n-(k-1)}H$ mod $I_{n-k}$, where $H\in S[X_{n},\ldots,X_{n-(k-1)}]$, $\deg(H)<d$ and $H$ is not a monomial. Then $\bar{F}_{k-1}=af_{e_{n-(k-1)}}\bar{H}$ mod $\bar{I}_{n-k}=0$ mod $\bar{I}_{n-k}$, since $\bar{F}=0$. This implies that $a\bar{H}\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Note that $\bar{I}_{n-k}=J_{G_{n-k}}$ and $e_{n-(k-1)}$ is a bridge in the graph $G_{n-k}\cup e_{n-(k-1)}$. It follows from Remark 1.2 that $a\bar{H}\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Lemma 4.1 it follows that $a\in J_{G_{n-k_{(e_{n-(k-1)})}}}:\bar{H}=J_{G_{n-k_{(e_{n-(k-1)})}}}$. Therefore, either $a\bar{H}=0$ or every monomial that survive in $a\bar{H}$ belongs to $J_{G_{n-k_{(e_{n-(k-1)})}}}$.
Subcase A. If $a\bar{H}=0$, then by the induction hypothesis, it follows that there is a linear relation, say $H^{\prime}=\sum_{i=1}^{n}b_{i}X_{i}$ such that $\bar{H^{\prime}}=0$, and $KH^{\prime}=aH$ for some $K\in R$. Then set $F_{k}=F_{k-1}-X_{n-(k-1)}KH^{\prime}$.
Subcase B. If $a\bar{H}\neq 0$ (i.e, there exists a monomial in the polynomial $aX_{n-(k-1)}H$ which survive on $af_{e_{n-(k-1)}}\bar{H}$), then choose leading term of $aX_{n-(k-1)}H$ with respect to any fixed term order which survive in $af_{e_{n-(k-1)}}\bar{H}$, say $cX_{n-(k-1)}^{\ell_{1}}\cdots X_{n}^{\ell_{k}}$, where $\lvert\ell_{1}\rvert+\cdots+\lvert\ell_{k}\rvert\leq d$ and $\ell_{1}>0$. We show that each such surviving monomial can be linearized. Then, the monomial $cf_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}$ belongs to $\bar{I}_{n-k}$, since $cf_{e_{n-(k-1)}}\in\bar{I}_{n-k}$ as $c\in J_{G_{n-k_{(e_{n-(k-1)})}}}$. Then from Proposition 2.1 and Definition 2.1 it follows that $c\in\bar{I}_{n-k}:f_{e_{n-(k-1)}}^{\ell_{1}}\cdots f_{e_{n}}^{\ell_{k}}=\bar{I}_{n-k}:f_{e_{n-(k-1)}}$. Therefore one has $cf_{e_{n-(k-1)}}-\sum_{i=1}^{n-k}b_{i}f_{e_{i}}=0$. Set $H_{k,1}=cX_{n-(k-1)}-\sum_{i=1}^{n-k}b_{i}X_{i}$ and $F_{k,1}=F_{k-1}-H_{k,1}(X_{n-(k-1)}^{\ell_{1}-1}\cdots X_{n}^{\ell_{k}})$. Repeat Case II for $F_{k,1}$. If it falls under either Case I or Subcase A, then proceed to step $k+1$. Otherwise, recursively define $H_{k,j}$ and $F_{k,j}$, and continue applying Case II to $F_{k,j}$ for $j\geq 2$, until it eventually falls under either Case I or Subcase A. This recursive process terminates in finitely many steps, as $aX_{n-(k-1)}H$ is a polynomial.
Case III. If $F_{k-1}(X_{1},\ldots,X_{n})=0$ mod $I_{n-k}$ then set $F_{k}=F_{k-1}$.
Observe that at each kth step, the claim reduces to showing that $F_{k}$ is of linear type. Hence, by the end of Step $n-1$, one can conclude that $F$ is of linear type. ∎
Corollary 4.1.
If $G$ be a tree then $J_{G}$ is a sequence of linear type.
Proof.
The proof follows from Theorem 3.1 and Theorem 4.1. ∎
In ([3, Corollary 6.2]), Bolognini et al. proved that if $G$ is a bipartite graph, then the binomial edge ideal $J_{G}$ coincides with the parity binomial edge ideal $\mathcal{I}_{G}$ (cf. [12]) and Lovász–Saks–Schrijver ideal $L_{G}(2)$ (cf. [8]), use which we establish the linear type properties of these ideals for trees in the following remark.
Remark 4.1.
By using [3, Corollary 6.2], one can obtain that if $G$ be a tree, then the associated parity binomial edge ideal and Lovász–Saks–Schrijver ideal are of linear type.
Corollary 4.2.
Suppose the sequence of elements $z_{1},\ldots,z_{n}$ in $S$ form a $p$-sequence such that it satisfies that the ideal $(z_{1},\ldots,z_{i})$ is prime for all $1\leq i\leq n$. Then the ideal $I=(z_{1},\ldots,z_{n})$ is of linear type.
4.1. Defining ideal of Rees algebra
The subsequent result follows from Theorem 3.1 and [11, Theorem 3.2], where we obtain a minimal generating set for the defining ideal of $\mathcal{R}(J_{G})$ when $G$ is a tree.
Corollary 4.3.
Let $\sigma:S[X_{\{i,j\}}\mid\{i,j\}\in E(G)]\rightarrow\mathcal{R}(J_{G})$ be the map defined by $\sigma(X_{\{i,j\}})=f_{ij}t$. If $G$ is a tree then the kernel of $\sigma$ (the defining ideal of $\mathcal{R}(J_{G})$) is minimally generated by { $f_{ij}X_{\{k,l\}}-f_{kl}X_{\{i,j\}}\mid\{i,j\}\neq\{k,l\}\in E(G)\cup$ $(-1)^{p_{A}(j)}f_{kl}X_{\{i,j\}}+(-1)^{p_{A}(k)}f_{jl}X_{\{i,l\}}+(-1)^{p_{A}(l)}f_{jk}X_{\{i,l\}}$, where $A=\{i,j,k,l\}$ is claw $(K_{1,3})$ with center at $i$ and $p_{A}(i)=\lvert\{j\in A\mid j\leq i\}\rvert$.
5. Counterexample
In this section, we present a counterexample that refutes Conjecture A specifically for a certain family of unicyclic graphs. Initially, we define the notation necessary for our demonstration.
Notation 5.1.
We denote by $C_{n,k}$ the class of graphs obtained by attaching $k$ pendant edges to each vertex of the cycle graph $C_{n}$.
Example 7.
[…]
Proposition 5.1.
If $C_{4,1}$ is an induced subgraph of $G$, then $G$ is not a linear type.
Proof.
Let $\phi:S[X_{ij}\mid\{i,j\}\in E(G)]\rightarrow S$ be a map given by $\phi(X_{ij})=f_{ij}$. Let $\Lambda$ be a kernel of $\phi$ (see, Definition 1.3). Suppose $C_{4,1}$ is an induced subgraph of $G$. Without loss of generality, we may assume that $C_{4,1}$ has the same vertex set and edge set as given in Example 7. Then, one can verify that
$\begin{split}&(x_{4}x_{8}y_{3}y_{7}-x_{3}x_{8}y_{4}y_{7}-x_{4}x_{7}y_{3}y_{8}+x_{3}x_{7}y_{4}y_{8})f_{15}f_{26}\\ &+(x_{6}x_{8}y_{4}y_{7}-x_{4}x_{8}y_{6}y_{7}-x_{6}x_{7}y_{4}y_{8}+x_{4}x_{7}y_{6}y_{8})f_{15}f_{23}\\ &+(x_{3}x_{8}y_{5}y_{7}-x_{5}x_{8}y_{3}y_{7}-x_{3}x_{7}y_{5}y_{8}+x_{5}x_{7}y_{3}y_{8})f_{14}f_{26}\\ &+(x_{5}x_{8}y_{6}y_{7}-x_{6}x_{8}y_{5}y_{7}-x_{5}x_{7}y_{6}y_{8}+x_{6}x_{7}y_{5}y_{8})f_{14}f_{23}\\ &+(x_{3}x_{7}y_{5}y_{6}-x_{3}x_{6}y_{5}y_{7}-x_{3}x_{5}y_{6}y_{7}+x_{5}x_{6}y_{3}y_{7})f_{12}f_{48}\\ &+(x_{4}x_{6}y_{5}y_{8}-x_{4}x_{8}y_{5}y_{6}-x_{5}x_{6}y_{4}y_{8}+x_{4}x_{5}y_{6}y_{8})f_{12}f_{37}\\ &+(x_{7}x_{8}y_{5}y_{6}-x_{6}x_{8}y_{5}y_{7}-x_{5}x_{7}y_{6}y_{8}+x_{5}x_{6}y_{7}y_{8})f_{12}f_{34}=0.\end{split}$ (5.1)
Therefore, one has
$\begin{split}F&=(x_{4}x_{8}y_{3}y_{7}-x_{3}x_{8}y_{4}y_{7}-x_{4}x_{7}y_{3}y_{8}+x_{3}x_{7}y_{4}y_{8})X_{15}X_{26}\\ &+(x_{6}x_{8}y_{4}y_{7}-x_{4}x_{8}y_{6}y_{7}-x_{6}x_{7}y_{4}y_{8}+x_{4}x_{7}y_{6}y_{8})X_{15}X_{23}\\ &+(x_{3}x_{8}y_{5}y_{7}-x_{5}x_{8}y_{3}y_{7}-x_{3}x_{7}y_{5}y_{8}+x_{5}x_{7}y_{3}y_{8})X_{14}X_{26}\\ &+(x_{5}x_{8}y_{6}y_{7}-x_{6}x_{8}y_{5}y_{7}-x_{5}x_{7}y_{6}y_{8}+x_{6}x_{7}y_{5}y_{8})X_{14}X_{23}\\ &+(x_{3}x_{7}y_{5}y_{6}-x_{3}x_{6}y_{5}y_{7}-x_{3}x_{5}y_{6}y_{7}+x_{5}x_{6}y_{3}y_{7})X_{12}X_{48}\\ &+(x_{4}x_{6}y_{5}y_{8}-x_{4}x_{8}y_{5}y_{6}-x_{5}x_{6}y_{4}y_{8}+x_{4}x_{5}y_{6}y_{8})X_{12}X_{37}\\ &+(x_{7}x_{8}y_{5}y_{6}-x_{6}x_{8}y_{5}y_{7}-x_{5}x_{7}y_{6}y_{8}+x_{5}x_{6}y_{7}y_{8})X_{12}X_{34}\in\Lambda_{2}\subset\Lambda.\end{split}$
From Remark 1.4, it follows that there is no linear relation in the first syzygy of $J_{G}$. This implies that $F$ is not an element of the module defined by the first syzygy of $J_{G}$. Therefore, $J_{G}$ is not a linear type, since $F$ is quadratic homogeneous element. ∎
Remark 5.1.
In particular, if a unicyclic graph $G$ contains $C_{4,1}$ as an induced subgraph, then $J_{G}$ is not of linear type.
Using Macaulay2 [6], we computed that if a graph $G$ contains $C_{6,1}$ as an induced subgraph, then $J_{G}$ is not of $2$ relation type. Thus, one can pose the following question in general:
Question 5.1.
Let $G$ be a graph. If $G$ contains $C_{2n,1}$ as an induced subgraph, then $J_{G}$ is not of linear type for all $n\geq 2$. In particular, $J_{G}$ is not of $n-1$ relation type.
Appendix
footnotetext: The appendix includes detailed verification of some of the computations presented in this manuscript. It may be omitted from the final version upon the referee’s suggestions
In this section, we present Macaulay2 [6] computations that verify the calculations done in the preceding sections.
Computation 5.1.
Verification of non-$p$-sequence property in Example 3.
1A=QQ[x_1..x_6]
2-- Input minimal generators of an ideal I.
3I={x_1*x_3*x_4*x_5,x_1^2*x_2*x_6,x_1^2*x_2^2*x_3*x_5}
4numGens=#I;
5-- Generate all permutations of the sequence
6l=permutations {x_1*x_3*x_4*x_5,x_1^2*x_2*x_6,x_1^2*x_2^2*x_3*x_5};
7numPerms=#l;
8
9-- Check \$p\$-sequence conditions for each permutation. If at least one false appears, then the sequence is not a \$p\$-sequence.
10
[…]
Computation 5.2.
Computation to verify Equation (5.1).
1S=QQ[x_1..x_8,y_1..y_8]
2-- Define the ideal X with the given generators
3X=ideal(x_1*y_2-x_2*y_1,x_2*y_3-x_3*y_2,x_3*y_4-x_4*y_3,x_1*y_4-x_4*y_1,x_1*y_5-x_5*y_1,x_6*y_2-x_2*y_6,x_7*y_3-x_3*y_7,x_8*y_4-x_4*y_8)
4
5-- Checking for linear type
6isLinearType X
7
8-- Compute the Rees ideal of X
9I=reesIdeal X
10
11
12-- Verify that the evaluation on the element I_35 yields 0
13
14(((x_4*x_8*y_3*y_7)-(x_3*x_8*y_4*y_7)-(x_4*x_7*y_3*y_8)+(x_3*x_7*y_4*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_5+x_5*y_1))+(((x_6*x_8*y_4*y_7)-(x_4*x_8*y_6*y_7)-(x_6*x_7*y_4*y_8)+(x_4*x_7*y_6*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_5+x_5*y_1))+(((x_3*x_8*y_5*y_7)-(x_5*x_8*y_3*y_7)-(x_3*x_7*y_5*y_8)+(x_5*x_7*y_3*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_4+x_4*y_1))+(((x_5*x_8*y_6*y_7)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_6*x_7*y_5*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_4+x_4*y_1))+(((x_3*x_7*y_5*y_6)-(x_3*x_6*y_5*y_7)-(x_3*x_5*y_6*y_7)+(x_5*x_6*y_3*y_7))*(-x_1*y_2+x_2*y_1)*(x_8*y_4-x_4*y_8))+(((x_4*x_6*y_5*y_8)-(x_4*x_8*y_5*y_6)-(x_5*x_6*y_4*y_8)+(x_4*x_5*y_6*y_8))*(x_7*y_3-x_3*y_7)*(-x_1*y_2+x_2*y_1))+(((x_7*x_8*y_5*y_6)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_5*x_6*y_7*y_8))*(-x_3*y_4+x_4*y_3)*(-x_1*y_2+x_2*y_1))
Acknowledgement. The first author thanks A. V. Jayanthan for his valuable comments. The first author thank the Scientific and Technological Research Council of Turkey - TÜBİTAK (Grant No: 124F113) for the financial support. The second author gratefully acknowledges partial funding received from the Science and Engineering Research Board (SERB), ANFR, India, under the Core Research Grant (CRG/2023/007668).
Competing Interests. The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.